OpenStudy (anonymous):

What is 2/3x = -1 4/5 ??

OpenStudy (anonymous):

ok leave x alone

OpenStudy (anonymous):

First you have to get rid of the fraction so you have to find the LCD.

OpenStudy (anonymous):

think -1 4/5 as a mixed fraction which is -9/5, now\[\frac{ 2 }{ 3 }x=\frac{ -9 }{ 5 }=>x=\frac{ -9 }{ 5 }*\frac{ 3 }{ 2 }=?\]

OpenStudy (anonymous):

1. Make -1 4/5 into a mixed fraction 2.You put the whole thing in parenthesis. (2/3x=-1 4/5) 3. Multiply it by the LCD 4. Then you have a 2 step equation